# SAQR-QC: A Logic for Scalable but Approximate Quantitative Reasoning about Quantum Circuits

### v1

@@ 4.3. First Attempt at Defining Judgments
Given a fixed domain $S=(s_{1},\cdots,s_{m})$ and a quantum program $\mathbf{C}=\lambda\rho.U\rho U^{\dagger}$, operating on a density matrix, we say that $\mathscr{A}=(A_{s_{1}},\ldots,A_{s_{m}})$ is a local precondition of $\mathscr{B}=(B_{s_{1}},\ldots,B_{s_{m}})$ if
(7) ${\mathrm{Tr}}\left(M_{\mathscr{A}}\rho\right)\leq{\mathrm{Tr}}\left(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)\right)$
for input state $\rho$, where $M_{\mathscr{A}}=\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ and $M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}$, which would correspond to the (candidate) judgment
$\{M_{\mathscr{A}}\}\ C\ \{M_{\mathscr{B}}\}.$
@@
where the inequality follows from 2.4 applied to the premise of Rule 8, and the first equality uses 2.2. Equation 9 gives us Equation 7, and thus 4.5 tells us that the conclusion of Rule 8 holds.
This definition naturally gives rise to a correctness judgment of the form $\{\mathscr{A}\}\ \mathbf{C}\ \{\mathscr{B}\}$ for quantum circuits. In addition to enabling backward reasoning, Rule 8 with 4.5 also supports forward reasoning by changing the premise of Rule 8 to
@@
Unfortunately, it is not possible, in general, to use either version of Rule 8 algorithmically to compute the weakest precondition and the strongest postcondition, respectively. The issue is one of expressibility: can the answer be decomposed according to the chosen scheme $S=(s_{1},\ldots,s_{m})$? In general, this structural constraint of our candidate logic presents an obstacle. For instance, even when $U_{F}$ is a two-qubit unitary, the transformed observable
@@
may not admit a decomposition of the form $\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ for any choice of local observables $A_{s_{i}}$. That is, the equality
@@
may have no solution for local predicates $\mathscr{A}$, even if $\mathscr{B}$ is given.
To see this limitation concretely, consider the following simple example where the circuit input is fixed as |$\ket{0}\ket{0}$.
@@
The program is illustrated as the following CNOT circuit
Figure 1. CNOT circuit $\mathbf{C}$
Let us reason about circuit $\mathbf{C}$ using Equation 10. We choose $(s_{1},s_{2})=(\{q_{1}\},\{q_{2}\})$ and $M_{\mathscr{B}}=|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$.
$U_{\mathbf{C}}^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U_{\mathbf{C}}=|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|.$
We now show that there is no $A_{1}\geq 0$ and $A_{2}\geq 0$ that satisfies
(11) $|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|=A_{1}\otimes I+I\otimes A_{2}.$
@@
{mdframed}
Observation: This calculation demonstrates that, in local reasoning, it is not always possible to express the weakest precondition.
We now show that under the QHL-like proof rule 8, sometimes, cannot prove some simple assertion. For instance, for the input |$\ket{00}$, the CNOT gate has no effect; therefore, the output state is also |$\ket{00}$. Thus, the following assertion holds:
@@
To see this, we compute
@@
[…]
@@
However, the final inequality is not valid. Therefore, according to Rule 8, Assertion 12 is not provable
⊬$\not\vdash\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}.$
@@ 4.4. Quantitative Judgments and Validity
In addition to the expressivity issues discussed in § 4.3, the predicates used in that section are not even capable of expressing the singleton quantum state |$\ket{0}^{\otimes n}$. This situation indicates that there is a mismatch with our requirements: 4.5 requires a triple to be valid for any state that satisfies the precondition predicate, whereas quantum programs typically focus on the specific input state |$\ket{0}^{\otimes n}$. For both of these reasons, we seek a more-expressive logic that still allows quantitative local reasoning.
Rather than using the predicate in 4.4 alone, we use quantum abstract interpretation (YP21, ) as an aid.
@@
for program $\mathbf{C}$ general predicates $\mathscr{A}$, $\mathscr{B}$ and projective predicates $\mathcal{P}$, $\mathcal{Q}$.
@@
A judgment $\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}$ as defined in 4.6 is true if we have $\forall\rho\vDash^{QAI}\mathcal{P}$,
@@
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are the matrix representations of $\mathscr{A}$ and $\mathscr{B}$, as defined in 4.2.
@@
in which case this definition simplifies to 4.5.
We have that ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ only depends on the reduced density matrices of $\rho$ with respect to the systems $\set{s_{i}}$, as we know from Equation 6,
@@
where the last step is due to 2.2.
Let $\mathscr{A}^{\prime}=(P_{s_{1}}A_{s_{1}}P_{s_{1}},\cdots,P_{s_{m}}A_{s_{m}}P_{s_{m}})$; we have
${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)={{\mathrm{Tr}}}(M_{\mathscr{A}^{\prime}}\rho).$
That is,
$\{\mathscr{A}^{\prime}|\mathcal{P}\}\mathbf{skip}\{\mathscr{A}|\mathcal{P}\}\ \ \&\ \ \{\mathscr{A}|\mathcal{P}\}\mathbf{skip}\{\mathscr{A}^{\prime}|\mathcal{P}\}$
In other words, the qualitative insights from quantum abstract interpretation enhance the precision of the reasoning by providing “stronger” postconditions or “weaker” preconditions.555 Although it is not clear the order relationship between $\mathscr{A}$ and $\mathscr{A}^{\prime}$. At least we have ${\mathrm{Tr}}(A_{s_{i}})\geq{\mathrm{Tr}}((P_{s_{i}}P_{s_{i}})A_{s_{i}})\geq{\mathrm{Tr}}(P_{s_{i}}A_{s_{i}}P_{s_{i}})$, where we used Lemma 2.2.
@@ 4.5. Reduction
We now consider the relationship between the correctness $\models^{\rm QAI}$ in QAI and $\models$ in 4.7.
@@
If $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ in the sense of 4.7, then $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$.
Proof.
Assume that $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ as defined in 4.7. Then for any $\rho$, we have
Let us choose $\rho\models\mathcal{P}$. According to Lemma 2.5, we have
On the other hand,
${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$
$=$ $\sum_{i}1=m.$
where the inequality follows from 2.4 and $Q_{s_{i}}\leq I_{s_{i}}$.
Therefore, ${\mathrm{Tr}}(M_{\mathscr{Q}}\llbracket\mathbf{C}\rrbracket(\rho))$ and ${\mathrm{Tr}}(Q_{s_{i}}{\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho))=1$ for all $i$. According to Lemma 2.5, we know ${\mathrm{Tr}}_{[n]\setminus s_{i}}\llbracket\mathbf{C}\rrbracket(\rho)\vDash Q_{s_{i}}$, which implies
$\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\},$
where Equation 3 $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$ iff for all $\rho$, $\rho\vDash\mathcal{P}$ implies $\llbracket\mathbf{C}\rrbracket(\rho)\vDash\mathcal{Q}$. ∎
@@ 4.6. Logical System with Soundness
The inference rules for program constructs in SAQR-QC are given in Figure 2.
@@
The proof system in Figure 2 is sound. That is, for quantum program $\mathbf{C}$, $\vdash\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}\ {\rm implies}\ \models\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}.$
@@
4.7. How to use the Rules for Scalable Reasoning
Among the proof rules in Figure 2, only the rule Unit presents a potential challenge to scalable reasoning. Specifically, the condition
@@ 4.7. How to use the Rules for Scalable Reasoning
requires computing the projector $\gamma(\mathcal{P})$, which is generally intractable for systems with a large number of qubits. That is, given a postcondition $\mathscr{B}$ along with projective predicates $\mathcal{P}$—it remains unclear how to compute a suitable precondition $\mathscr{A}$ in a scalable manner.
To address this issue, we must allow for controlled approximations that trade some precision for tractability. In this subsection, we propose methods for applying the rule Unit while preserving scalability. The core idea is to construct inequalities over high-dimensional systems by composing inequalities over smaller, lower-dimensional subsystems. This compositional approach ensures that reasoning remains feasible, even as the number of qubits in the quantum system increases.
Warm-up exercise: Ignore QAI: As a warm-up exercise—which is potentially subject to the same kind of imprecision that we saw in § 4.3), but sometimes gives a precise-enough result (as we will see in § 5)—we can use the following inequality, in which we ignore all information from the QAI component (i.e., $\gamma(\mathcal{P})$):
@@
(16) $M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\Longleftrightarrow\sum_{i}{A_{s_{i}}}\otimes I_{[n]\setminus s_{i}}\leq U^{{\dagger}}\sum_{i}{B_{s_{i}}}\otimes I_{[n]\setminus s_{i}}U.$
This inequality is linear, so it can be satisfied by fulfilling a set of inequalities in some number of smaller systems (indexed by $j$). We partition the domain of predicates $S=(s_{1},\ldots,s_{m})$, where each $s_{i}\subseteq[n]$. For $\cup_{j}T_{j}=\{1,2,\cdots,m\}$ with $T_{i}\cap T_{k}=\emptyset$ for all $i\neq k$, we can deal with $j$ inequalities as follows:
@@
(17) $\textrm{for all $j$}\qquad\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\leq U^{{\dagger}}\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}U\Longrightarrow M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}}).$
Alternatively, for forward reasoning, Equation 17 takes the form
(18) $\textrm{for all $j$}\qquad U\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}U^{{\dagger}}\leq\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\Longrightarrow M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}}).$
A more precise approach: Ignoring $\gamma(\mathcal{P})$ away may lose too much precision. We can refine the approach of the warm-up exercise as follows. We again consider some number of smaller systems (indexed by $j$) by partitioning the domain of predicates: for $\cup_{j}T_{j}=\{1,2,\cdots,m\}$ with $T_{i}\cap T_{k}=\emptyset$ for all $i\neq k$. Clearly, we have
(19) $\gamma(\mathcal{P})=\cap_{i}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}\subseteq P_{j}::=\cap_{i\in T_{j}}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Now we can use the following inequalities in our reasoning steps:
(20) for all $\textrm{for all $j$}\qquad P_{j}\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)P_{j}\leq P_{j}U^{{\dagger}}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)UP_{j}.$
Thus, our two basic reasoning tools will be Equation 20 along with Equations 17 and 18.
Equation 20 implies Equation 15. The proof is given in Appendix C.

### v3

@@ 3.2. A First Attempt at Defining Judgments
Given a fixed domain $S=(s_{1},\dots,s_{m})$ and a quantum program $\mathbf{C}=\lambda\rho.\,U\rho U^{\dagger}$ on a density matrix, we call $\mathscr{A}=(A_{s_{1}},\dots,A_{s_{m}})$ a local precondition of $\mathscr{B}=(B_{s_{1}},\dots,B_{s_{m}})$ if
$\{M_{\mathscr{A}}\}\,C\,\{M_{\mathscr{B}}\},$
where such a judgment means
$\models\{M_{\mathscr{A}}\}\,C\,\{M_{\mathscr{B}}\}\quad\mathrm{iff}\quad\textrm{for all}~\rho,\,~{\mathrm{Tr}}(M_{\mathscr{A}}\rho)\leq{\mathrm{Tr}}(M_{\mathscr{B}}\llbracket\mathbf{C}\rrbracket(\rho)),$
with $M_{\mathscr{A}}=\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ and $M_{\mathscr{B}}=\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}$.
@@
where the inequality follows from ˜2.4 applied to the premise of Rule 5, and the first equality uses ˜2.2. ˜3.6 tells us that the conclusion of Rule 5 holds.
This definition naturally gives rise to a correctness judgment of the form $\{\mathscr{A}\}\ \mathbf{C}\ \{\mathscr{B}\}$ for quantum circuits. In addition to enabling backward reasoning, Rule 5 with ˜3.6 also supports forward reasoning by changing the premise of Rule 5 to
@@
Unfortunately, it is not possible, in general, to use either version of Rule 5 algorithmically to compute the weakest precondition and the strongest postcondition, respectively. The issue is one of expressibility: can the answer be decomposed according to the chosen scheme $S=(s_{1},\ldots,s_{m})$? In general, this structural constraint of our candidate logic presents an obstacle. For instance, even when $U_{F}$ is a two-qubit unitary, the transformed observable
@@
may not admit a decomposition of the form $\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}$ for any choice of local observables $A_{s_{i}}$. That is, the following equality may have no solution for local predicates $\mathscr{A}$, for some given $\mathscr{B}$.
@@
To understand this limitation concretely, consider a circuit in which a CNOT gate is applied to the input state $|0\ket$$\ket{0}$, as shown in Figure 1, and let us reason about the circuit $\mathbf{C}$ using Equation˜7. We choose $(s_{1},s_{2})=(\{q_{1}\},\{q_{2}\})$ and define the postcondition to be $M_{\mathscr{B}}=|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$. Observation: The following calculation will demonstrate that, in local reasoning, it is not always possible to express the weakest precondition.
We will show that there is no $A_{1},A_{2}$ such that
$U_{\mathbf{C}}^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U_{\mathbf{C}}$ $=|0\rangle\langle 0|\otimes I+|0\rangle\langle 0|\otimes|0\rangle\langle 0|+|1\rangle\langle 1|\otimes|1\rangle\langle 1|$
$=A_{1}\otimes I+I\otimes A_{2}$
Figure 1. CNOT circuit $\mathbf{C}$.
@@
[…]
@@
To establish that Rule 5 cannot prove that Assertion 8 holds, we compute
@@
$\Longleftrightarrow$ $\mathit{CNOT}(|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|)\mathit{CNOT}^{{\dagger}}\leq|0\rangle\langle 0|\otimes I+I\otimes|0\rangle\langle 0|$
@@
Because the final inequality fails, in our strawman logic based on tuples of local observables and Rule 5, Assertion 8 is unprovable; i.e.,
$\not\vdash\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}\mathbf{C}\{(|0\rangle\langle 0|,|0\rangle\langle 0|)\}.$
@@ 3.3. Quantitative Judgments and Validity
[…]
@@
for program $\mathbf{C}$, general predicates $\mathscr{A}$ and $\mathscr{B}$, and projective predicates $\mathcal{P}$ and $\mathcal{Q}$.
@@
The judgment $\{\mathscr{A}|\mathcal{P}\}\,\mathbf{C}\,\{\mathscr{B}|\mathcal{Q}\}$ in ˜3.7 holds if $\forall\rho\vDash^{QAI}\mathcal{P}$,
@@
where $M_{\mathscr{A}}$ and $M_{\mathscr{B}}$ are the matrix representations of $\mathscr{A}$ and $\mathscr{B}$, as defined in ˜$3.$
@@
in which case this definition simplifies to ˜3.6.
We have that ${{\mathrm{Tr}}}(M_{\mathscr{A}}\rho)$ only depends on the reduced density matrices of $\rho$ with respect to the systems $\set{s_{i}}$, as we know from Equation˜4,
@@
where the last step is due to ˜$2.$
We observe that ${\mathrm{Tr}}(M_{\mathscr{A}}\rho)={\mathrm{Tr}}(M_{\mathscr{A}^{\prime}}\rho)$ with $\mathscr{A}^{\prime}=(P_{s_{1}}A_{s_{1}}P_{s_{1}},\dots,P_{s_{m}}A_{s_{m}}P_{s_{m}})$. Then
$\{\mathscr{A}^{\prime}|\mathcal{P}\}\,\mathbf{skip}\,\{\mathscr{A}|\mathcal{P}\}\quad\text{and}\quad\{\mathscr{A}|\mathcal{P}\}\,\mathbf{skip}\,\{\mathscr{A}^{\prime}|\mathcal{P}\}.$
Thus, the qualitative insights from quantum abstract interpretation refine the reasoning process by yielding “stronger" postconditions and “weaker" preconditions. [^2] Here, “stronger” and “weaker” simply refer to the relation $\mathrm{tr}(A_{s_{i}})\geq\mathrm{tr}\big((P_{s_{i}}P_{s_{i}})A_{s_{i}}\big)\geq\mathrm{tr}\big(P_{s_{i}}A_{s_{i}}P_{s_{i}}\big),$ as established by Lemma 2.2. They do not indicate any ordering relationship between $\mathscr{A}$ and $\mathscr{A}^{\prime}$.
@@ 3.4. Reduction
We now consider the relationship between the correctness $\models^{\rm QAI}$ in QAI and $\models$ in ˜3.8.
@@
If $\models\{\mathscr{P}\}C\{\mathscr{Q}\}$ in the sense of ˜3.8, then $\models^{\rm QAI}\{\mathcal{P}\}C\{\mathcal{Q}\}$.
We defer the proof to the appendix.
@@ 3.5. Logical System with Soundness
The inference rules for program constructs in SAQR-QC are given in Figure˜2.
@@
The proof system in Figure˜2 is sound. That is, for quantum program $\mathbf{C}$, $\vdash\{\mathscr{A}\}\mathbf{C}\{\mathscr{B}\}\ {\rm implies}\ \models\{\mathscr{A}|\mathcal{P}\}\mathbf{C}\{\mathscr{B}|\mathcal{Q}\}.$
@@
3.6. Scalable Application of the Unit Rule
Among the proof rules in Figure 2, only the rule Unit poses a significant challenge to scalability. Specifically, the condition
@@ 3.6. Scalable Application of the Unit Rule
requires computing the projector $\gamma(\mathcal{P})$, which becomes intractable for systems with many qubits. That is, given a postcondition $\mathscr{B}$ and a set of projective predicates $\mathcal{P}$, there is no scalable method for synthesizing a precondition $\mathscr{A}$ satisfying this inequality.
To address this issue, we introduce a compositional approximation strategy. The key idea is to construct inequalities over high-dimensional systems by composing inequalities over smaller subsystems. These approximations trade some precision for tractability, but preserve overall soundness and enable reasoning about large quantum programs.
Warm-up: Ignoring $\gamma(\mathcal{P})$
As a first step—which, although imprecise (see §˜3.2), may still yield useful results (as in §˜4)—we omit the projection $\gamma(\mathcal{P})$. This approximation gives rise to a simple linear inequality:
@@
(12) $M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}})\quad\Longleftrightarrow\quad\sum_{i}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\leq U^{\dagger}\left(\sum_{i}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U.$
To improve scalability, we partition the index set $\{1,\dots,m\}$ into disjoint subsets $T_{1},\dots,T_{k}$, where each $T_{j}$ corresponds to a group of local predicates. We establish soundness via localized reasoning:
@@
(13) $\forall j.\quad\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\leq U^{\dagger}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U\quad\Longrightarrow\quad M_{\mathscr{A}}\leq\llbracket\mathbf{C}\rrbracket^{*}(M_{\mathscr{B}}).$
The forward form of this inequality is:
(14) $\forall j.\quad U\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)U^{\dagger}\leq\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
Because each inequality involves a small subsystem, constraints can be efficiently discharged using symbolic solvers or semidefinite programming, depending on the structure of $A_{s_{i}}$ and $B_{s_{i}}$.
A More Precise Approach: Guarded Inequalities
The warm-up approach may lose critical precision due to ignoring $\gamma(\mathcal{P})$. To refine this, we incorporate the projection information in a compositional way. Let $\mathcal{P}=\{P_{s_{1}},\dots,P_{s_{m}}\}$, and again partition the index set into disjoint subsets $T_{1},\dots,T_{k}$. For each $T_{j}$, we define a local over-approximation of the global projector:
(15) $\gamma(\mathcal{P})\subseteq P_{j}:=\bigcap_{i\in T_{j}}P_{s_{i}}\otimes I_{[n]\setminus s_{i}}.$
We can write guarded versions of the local inequalities:
(16) $\forall j.\quad P_{j}\left(\sum_{i\in T_{j}}A_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)P_{j}\leq P_{j}U^{\dagger}\left(\sum_{i\in T_{j}}B_{s_{i}}\otimes I_{[n]\setminus s_{i}}\right)UP_{j}.$
These inequalities are more precise than those in (13), as they account for the structure imposed by $\mathcal{P}$, while still avoiding the need to compute the full projector $\gamma(\mathcal{P})$.
Summary of Reasoning Tools
We thus obtain two practical tools for scalable reasoning under the Unit rule:
  • Unguarded inequalities: (13) and (14), which enable fast reasoning but may be imprecise;
  • Guarded inequalities: (16), which retain more logical structure from $\gamma(\mathcal{P})$ while remaining tractable.
Finally, in Appendix˜D, we prove that the guarded inequalities imply the full condition (11), thereby validating the correctness of our compositional approximation framework.
Remark: The methodology presented in this section is general. For a given circuit, useful predicates can typically be constructed straightforwardly because each gate in a quantum circuit is represented by a local unitary operator that typically acts on at most two qubits. In each reasoning step, we analyze a single constant-qubit unitary—typically a one- or two-qubit gate—and extract the predicate $B_{s_{i}}$ over its qubits, leaving the remainder untouched. This approach is sound because the unitary does not affect the reduced density matrices of qubits outside its support. The convenience and effectiveness of this approach are illustrated in the examples throughout the paper.